# Compositional Hardness of Code in Large Language Models - A Probabilistic Perspective

### v1

@@ 1 Introduction
Large language models (LLMs), based on the transformer archietecture, Vaswani et al. (2017), have become very efficient problem solvers in many domains, such as broad-scoped question answering, writing assistance, teaching, and more Brown (2020); Radford et al. (2019); OpenAI (2023); Bubeck et al. (2023); Nori et al. (2023); West (2023). Yet their analytical skills, such as coding capabilities, are slow to develop - Chen et al. (2021b); Li et al. (2022a); Alp (2023); Ridnik et al. (2024) show that even with millions of generations, LLMs may not produce a single correct solution to competitive coding problems. Dziri et al. (2024) show that LLMs solve compositional tasks such as long multiplication and dynamic programming without developing systematic problem-solving skills.
@@
One way to empower LLMs in analytical tasks, is to use subtask decomposition, otherwise known as chain of thought (COT) - a method in which an LLM breaks down a problem to smaller, more manageable tasks, solves them, and integrates it into a solution. The method has been empirically demonstrated by Wei et al. (2022), that show reasoning capabilities of language models improve when they are prompted to break down a task. Its efficiency has also been studied theoretically - Wies et al. (2022); Malach (2023) prove that through the autoregressive nature of language models, problems that cannot be solved directly, can be solved by subtask decomposition, Merrill & Sabharwal (2023) prove that while transformers are limited in the computational problems they can solve directly, using a polynomial number of intermediate steps, they can represent any polynomial time Turing Machine, and Sanford et al. (2024) provide a similar result on expressing more complex arithmetic circuits using more steps of COT.
@@
Yet even with task decomposition, there is a limitation to transformer based models on analytical tasks with COT, due to their limited ability to compose functions - Sanford et al. (2024) show single layer attention can only learn pairwise relations between tokens, limiting the ability to integrate intermediate steps for a fixed model size, and the work of Peng et al. (2024) shows that iterative composition over a domain is limited by the size of the model, even when COT is used. Thus while theoretically possible, some tasks require an arbitrarily long COT for an LLM to solve. However, in practice, LLMs are limited in their context length - beyond the constraint of context length during training, Hsieh et al. (2024) introduce the RULER benchmark, for measuring the utility of LLMs on long context tasks. They show that in practice many models can perform tasks only on a much shorter context length than they were trained on. Similarly, Ebrahimi et al. (2024) show empirically that LLMs are limited in random access to tokens within the context, in the bit parity problem. Consequently, even though COT can in theory allow an LLM to solve arbitrarily complex analytical problems, in practice, they will be limited by the effective context length.
@@
A rising approach to remedy this limitation is to solve problems through the use of multi-agent systems, that tackle complex problems through the use of agents, where each agent is an LLM instance that solves a different aspect of the problem. While it has been used for simulating social interactions, Park et al. (2023); Li et al. (2023); Pang et al. (2024), it has also been shown as an effective tool for analytical problem solving. This can be done by decomposing a large task and distributing the sub-tasks between agents. Ishibashi & Nishimura (2024) use this method for building large code bases and Liu et al. (2023), use a dynamic LLM-agent network for solving code problems and analytical tasks.
@@
In this work, we theoretically study a compositional hardness of coding problems originating from context processing limitations of LLMs, and the resultant effectiveness of a multi-agent system over a single model instance in composite coding problems. We model a composite coding problem using a pair of simpler coding problems, where the output to the composite problem can be a basic manipulation on the outputs to the pair of problems (e.g. composition, product of outputs, concatenation, logical function of the outputs, etc.), such that the solution to the problem can be obtained from concatenating the solutions to the problem pair. The model’s usefulness on a coding task, is quantified by a generation complexity metric (definition 2) - the number of LLM generations required to sample at least one correct solution. The appeal of this metric, is that due to the existence of code testing units, it suffices to turn an LLM into a good program candidate generator and simply output a candidate that is correct Kambhampati et al. ; Thakur et al. (2023); Luo et al. (2024). The single LLM instance solves the entire problem, while the multi-agent system is comprised of two agents, each is an LLM instance tasked with solving one of the problems in the pair. Thus the LLM’s helpfulness in the single instance case is the generation complexity for the composite problem, while in the multi-agent case, it is the product of generation complexities of the pairs of problems, as a correct solution is attained when independently sampling a correct solution to each problem.
@@
We theoretically model an LLM as an autoregressive model, where a solution is sampled token by token, based on the hidden representations. When combining two problems whose solutions are grammatically similar but semantically different (such as different coding problems), we assume this combination injects noise into the model’s representations during solution generation for each sub-problem (assumption 1), which we denote as screening. We show that a compositional problem may have an exponential gap in generation complexity relative to the product of sub problems’ generation complexities (theorem 1), meaning an exponential hardness of composition in-context. Essentially, the model is less capable of solving two problems if they are presented within the same context, than if presented in separate contexts. This points to an advantage of decomposing a problem not only within the same LLM context, but to distribute the problems among multiple agents (i.e. solve each sub-task within a different context). Additionally, this result provides a view of the model’s effective context length through the lens of screening, which is the model’s ability to isolate the relevant context at each decoding step – additional irrelevant context may reduce model’s performance on other tasks within the context window exponentially with length. We validate our assumptions and results experimentally on Llama 3 by constructing composite code problems.
@@ 2.1 Generation Complexity
We focus on coding problems, meaning each problem is written in natural language and is solved by a function, and the goal is for the model to generate a code that implements it.
@@
We formally define the generation complexity of a program to a problem as the inverse success rate of a model to generate a correct solution to the problem:
@@
For a problem $x$ and a natural language distribution $\mathbb{P}$ over $\Sigma^{*}$, the generation complexity of $x$ w.r.t. $\mathbb{P}$, is:
$N(\mathbb{P},x)=\frac{1}{\sum_{y\in\textit{correct solutions}}\mathbb{P}(y|x)}$ (1)
@@
Intuitively, the generation complexity is the number of program candidates needed to be sampled from the distribution to get a program that solves the problem.
In this work, we consider simple compositional problems, which are decomposable to two smaller problems, and their concatenated solutions solve the compositional problem. This captures cases such as a composition of two code functions, or simple manipulations on outputs of code functions (e.g. concatenation, product, etc.). For a compositional problem $x$, decomposable to $x_{1},x_{2}$, we quantify the in-context hardness of composition as the ratio of generation complexity to the full problem $N(\mathbb{P},x)$ with the product of generation complexities for the sub-problems $N(\mathbb{P},x_{1})\cdot N(\mathbb{P},x_{2})$. If the latter is much smaller, this points to an advantage for distributing the sub-problems between different instances of an LLM (multiple agents), as it would require fewer generations to sample a correct solution. This defines a compositional hardness of coding problems that originates from context processing limitations, as the compositional problem can be mathematically equivalent to solving both problems, yet seeing them both in the same context reduces the model’s performance on them.
@@ 2.2 Screening in Autoregressive Models
Here we introduce a source of hardness in code composition based on the autoregressive nature of LLMs. Typically, latent representations of the model contain information about the context beyond the next token prediction, e.g. the structure of the solution to problems Ye et al. (2024). Thus when composing two code problems, we expect the representations during the generation of the second program to contain information about the first program and vice versa, which is grammatically similar (same programming language) but semantically very different. As a result, this additional information creates noise that can harm the generation process.
Formally, we denote by $r^{(L)}(x)$ the model’s last hidden layer representation of the prompt $x$, by $U$, the model’s unembedding matrix (hidden dimension to vocabulary). The logit of the $i$’th token is thus defined as the token’s score before the softmax operation: $\langle r^{(L)}(x),U^{T}e_{i}\rangle$, where $e_{i}$ is the one-hot vector of the token. The probability distribution at each decoding step is the softmax applied to the logits, $\mathbb{P}_{LLM}(i|x)=softmax(\langle r^{(L)}(x),U^{T}e_{i}\rangle)$.
@@
Essentially, this means the model attempts to generate the same solutions as in the non-compositional case, but noise may interfere in the process. The projection of this noise onto the dictionary creates noise in the logits during decoding, which can lead the model to make mistakes.
As the two problems $x_{1},x_{2}$ and their solutions $y_{1},y_{2}$ may be different semantically, we do not expect the noise to "push" the model towards the correct solutions more than to incorrect solutions, thus when projected onto the vocabulary, $V$, the noise on the logit of the correct token minus the noise on an incorrect token, $\langle noise,U^{T}e_{i_{\textit{correct next token}}}\rangle-\langle noise,U^{T}e_{i_{\textit{an incorrect next token}}}\rangle$, should be symmetric on average. Additionally it should be bounded within some range $[-M,+M]$, as it changes the hidden representation to a finite extent. In practice, we only expect this to be true for the high probability tokens, as the vocabulary $V$ is very large, and some low probability tokens may be systematically enhanced. To avoid this issue, we make our assumptions only on the weighted average of the noise, where the weights are given by the probability mass that the model assigns them. This way, low probability tokens with asymmetric noise or large norms receive low weight and are averaged with the noise of other tokens. Denote by $\mathbb{P}(i|context)$ the probability assigned to the $i$’th token given the context. We will make our assumptions on the weighted average of the noise on the incorrect token logits minus correct token logits:
$X=\frac{\sum_{i\in V/\{\textit{correct next token}\}}\mathbb{P}(i|context)\langle noise,U^{T}e_{i}-U^{T}e_{i_{\textit{correct next token}}}\rangle}{\sum_{i\in V/\{\textit{correct next token}\}}\mathbb{P}(i|context)}$ (4)

### v3

@@ 1 Introduction
Large language models (LLMs), based on the transformer archietecture (Vaswani et al., 2017), have become very efficient problem solvers in many domains, such as broad-scoped question answering, writing assistance, teaching, and more (Brown, 2020; Radford et al., 2019; OpenAI, 2023; Bubeck et al., 2023; Nori et al., 2023; West, 2023). Yet their analytical skills, such as coding capabilities, were slow to develop - Chen et al. (2021b); Li et al. (2022a); Alp (2023); Ridnik et al. (2024) showed LLMs struggle on competitive coding problems, and  Zhuo et al. (2024) showed they struggle on complex coding tasks.
@@
One way to empower LLMs in analytical tasks, is to use subtask decomposition, otherwise known as chain of thought (CoT) – a method in which an LLM breaks down a problem to smaller, more manageable tasks, solves them, and integrates it into a solution. The method has been empirically demonstrated by Wei et al. (2022), that show reasoning capabilities of language models improve when they are prompted to break down a task. Since then, the method has been incorporated into the training phase of LLMs, in models such as openAI’s o1 Jaech et al. (2024), Deep-Seek-R1 Guo et al. (2025), QWQ Qwen (2024) and Kimi K1.5 Team et al. (2025), achieving state of the art results on existing coding benchmarks. Its efficiency has also been studied theoretically - Wies et al. (2022); Malach (2023); Merrill & Sabharwal (2023); Sanford et al. (2024) proving through the autoregressive nature of language models, that problems that cannot be solved directly, can be solved by subtask decomposition, allowing to represent any polynomial time turing machine with a polynomial number of CoT steps.
@@
Yet even with task decomposition, due to the transformer architecture’s limited ability to compose functions and learn complex relations between tokens Sanford et al. (2024); Peng et al. (2024); Xu et al. (2024), some tasks require an arbitrarily long CoT for an LLM to solve. This is especially prominant in SOTA LLMs such as openAI’s o3, that generates $\sim 100k$ tokens per problem on the ARC AGI benchmark Chollet (2024). However, in practice, LLMs are limited in their context length – beyond the constraint of context length during training, Hsieh et al. (2024); Liu et al. (2024) show that in practice many models can perform tasks only on a much shorter context length than they were trained on and cannot fully use all the information within their context. Consequently, even though CoT can in theory allow an LLM to solve arbitrarily complex analytical problems, in practice, they will be limited by the effective context length. This is especially concerning for more practical programming applications than those considered in coding benchmarks, such as building and interacting with large code bases, too big for any LLM context to contain.
@@
A rising approach to remedy this limitation is to solve problems through the use of multi-agent systems, that tackle complex problems through the use of agents, where each agent is an LLM instance that solves a different aspect of the problem. While it has been used for simulating social interactions, (Park et al., 2023; Li et al., 2023; Pang et al., 2024), it has also been shown as an effective tool for analytical problem solving. This can be done by decomposing a large task and distributing the sub-tasks between agents. Ishibashi & Nishimura (2024) use this method for building large code bases and Liu et al. (2023), use an LLM-agent network for solving code problems and analytical tasks.
@@
In this work, we theoretically study a compositional hardness of coding problems originating from context processing limitations of LLMs, and the resultant effectiveness of a multi-agent system over a single model instance in composite coding problems. We model a coding problem as a composition of simpler sub-problems, as is typically invoked by LLMs in a CoT process, such that combining the codes to all the sub-problems creates a program that implements the full solution. The model’s usefulness on a coding task is quantified by a generation complexity metric (definition 2) – the number of LLM generations required to sample at least one correct solution. The appeal of this metric is that due to the existence of code testing units, it suffices to turn an LLM into a good program candidate generator and simply output a candidate that is correct (Kambhampati et al., ; Thakur et al., 2023; Luo et al., 2024). When a single LLM instance solves all the sub-problems in a single context, its usefulness is the generation complexity for the composite problem. The multi-agent on the other hand, distributes the sub-problems between LLM instances, sampling a corresponding solution from each, its helpfulness is the product of generation complexities of the sub-problems, corresponding to sampling independently a correct solution to each.
@@
We theoretically model an LLM as an autoregressive model, where a solution is sampled token by token, based on the hidden representations. When combining two problems whose solutions are grammatically similar but semantically different (such as different coding problems) into a single context window, we assume this combination injects noise into the model’s representations during solution generation for each sub-problem (assumption 1), which we denote as screening. We show that a compositional problem may have an exponential gap in generation complexity relative to the product of sub problems’ generation complexities (theorem 1), meaning an exponential hardness of composition in-context. Essentially, the model is less capable of solving two problems if they are presented within the same context, than if presented in separate contexts. This points to an advantage of decomposing a problem not only within the same LLM context, but to distribute the problems among multiple agents (i.e. solve each sub-task within a different context). Additionally, this result provides a view of the model’s effective context length through the lens of screening, which is the model’s ability to isolate the relevant context at each decoding step – additional irrelevant context may reduce model’s performance on other tasks within the context window exponentially with length. We validate our assumptions and results experimentally on Llama 3 8B Instruct and 70B Instruct, and Mistral 7B Instruct V0.3 by constructing composite code problems.
2 Related Works
LLMs as solution candidate generators in programming:
Past empirical works have shown that while LLMs struggle to solve complex programming tasks with a single sampled solution, on some problems sampling thousands of solutions may yield a handful of correct ones that can be detected by testing units (Chen et al., 2021b; Li et al., 2022a; Alp, 2023; Ridnik et al., 2024). Hence empowering an LLM with a large sampling budget drastically boosts its capabilities. This procedure has of recent gained popularity under the term "scaling test time compute", which aims at leveraging compute at inference time Snell et al. (2024), showing it may be more advantageous than using it for additional training time. This has been recently demonstrated at the extreme scale with the results of OpenAI’s o3 on the ARC AGI challenge Chollet (2024), where an increased budget of 1024 parallel samples per problem from the model contributed to a significantly higher performance than using 6 parallel samples. As such, a relevant approach to quantify an LLM’s usefulness in such tasks is the expected number of programs one needs to sample from the model to obtain a correct solution. This motivates our definition of generation complexity (definition 2), and in this work, we primarily focus on providing theoretical results on the relative generation complexity in compositional coding tasks.
Theoretical results on composition:
Previous works on multi-step function composition and subtask decomposition in language models primarily focus on expressing and learning a deterministic function that solves well defined mathematical tasks Ahuja & Mansouri (2024), such as solving the bit parity problem (Wies et al., 2022), expressing and learning Turing machines (Merrill & Sabharwal, 2023; Malach, 2023), composing a mathematical function arbitrary numbers of times (Peng et al., 2024), or simple mathematical operations on text Xu et al. (2024). These do not translate in a natural way to contemporary code generation methods using LLMs, which are based on probabilistic sampling in order to explore the vast solution space, as in the above works, the model provides a deterministic answer which is either correct or incorrect and the notion of sampling multiple solutions cannot be incorporated. Hence beyond the fact that SOTA LLMs do not deterministically generate solutions to coding problems, the importance of scaling test time compute is absent. In this work, we deal with probabilistic autoregressive generation of solutions to problems, which is more applicable to the practice of LLM code generation. We provide a softer and more informative result for how hard it is to compose problems with the generation complexity metric, which is not possible through previous approaches and incorporate the existence of multiple correct solutions to a single problem, which is typical in programming.
Effectiveness of LLMs in utilizng long context:
Previous empirical works (Hsieh et al., 2024; Liu et al., 2024) have shown LLM performance on tasks such as retrieval degrade when performed on longer contexts, and that models may be limited in their random access to tokens within the context (Ebrahimi et al., 2024). In this work, we study the degradation of language models on more complex tasks that involve not just retrieval operations from a given long context, but also include long text generation: Namely, compositional coding tasks through the lens of noise that pieces of context from different subtasks insert into the generation process. Differently from above mentioned works, in our results the context does not necessarily have to be very long in order for the model performance to drop, but rather that grammatically similar but semantically different pieces of the context may “confuse" the model and harm code generation even in shorter contexts.
@@ 3.1 Generation Complexity
We focus on coding problems, meaning each problem is written in natural language and is solved by a function, and the goal is for the model to generate a code implementing it.
@@
We formally define the generation complexity of a problem as the inverse success rate of a model to generate a correct solution to the problem:
@@
For a problem $x$ and a natural language distribution $P$ over $V^{*}$, the generation complexity of $x$ w.r.t. $P$, is:
$N(P,x)=\frac{1}{\sum_{y\in\textit{correct solutions}}P(y|x)}$ (1)
Where $V^{*}$ is the Kleene closure of the vocabulary $V$.
@@
Intuitively, the generation complexity is the number of program candidates needed to be sampled from the conditional distribution $P(\cdot|x)$ to get a program that solves the problem.
Emergence of compositional coding problems from general coding problems in LLMs with CoT:
In this work, we are interested in deriving generation complexity bounds for coding problems. Generally, LLMs solve coding problems by producing a CoT that suggests a decomposition of the problem $x$, to $k$ subproblems $x_{1}...x_{k}$. Each subtask is then implemented by the model as a block of code $y_{1}...y_{k}$, and the concatenation of these blocks $y_{1}\oplus...\oplus y_{k}$ creates a single program that solves the problem. We note that the different blocks can interact, hence generally it is a compositional problem with $k$ components. Thus all but the simplest coding problems are mapped by LLMs into compositional coding problems of the form described above.
Decoupling hardness of decomposition from hardness of composition in general problems:
Based on the above, the success rate of a model to generate a correct solution to a problem is determined by two factors. The first is decomposition – generating a CoT $x_{1}…x_{k}$ with a correct decomposition to the problem. The second is composition – to implement a program $y_{1}\oplus…\oplus y_{k}$ matching the subtasks described in the CoT. In the following, we isolate the contribution of composition to generation complexity and show it bounds the generation complexity to the full problem.
Denote the set of valid decompositions to the problem $x$ as: $d(x)=\{x_{1},…,x_{k}|x_{1},…,x_{k}\textit{ is a valid decomposition to x}\}$ and the set of correct solutions to a subtask $x_{i}$ as: $s(x_{i})=\{y|\textit{y is a correct solution to $x_{i}$}\}$. The probability that a model generates a correct solution to the problem equals the probability to generate a correct decomposition $P(x_{1}...x_{k}|x)$, times the probability to implement a correct solution based on the decomposition, which in terms of generation complexity is $\frac{1}{N(P,(x,x_{1}...x_{k}))}$. We arrive at the following formula for generation complexity to a general problem as a weighted sum of compositional problems:
$\frac{1}{N(P,x)}=\sum_{(x_{1}...x_{k})\in d(x)}\frac{P(x_{1},...,x_{k}|x)}{N(P,(x,x_{1}...x_{k}))}$ (2)
We see the manifestation of hardness of searching for a decomposition in the numerator, and of implementing a correct solution to a compositional problem in the denominator. The compositional aspect is interesting since models are famously not good at ignoring irrelevant information, which can harm the ability to solve multiple sub-tasks within the same context. Hence even given a correct decomposition to the problem, implementing it could still prove a challenge. We thus lower bound the generation complexity to the problem in terms of the compositional hardness:
$N(P,x)\geq\min_{(x_{1}...x_{k})\in d(x):P(x_{1}...x_{k}|x)>0}N(P,(x,x_{1}...x_{k}))$ (3)
Meaning the generation complexity is at least as large as the generation complexity of solving the decomposed problem for the best decomposition that the model can suggest. In our main results, we find a non-trivial upper bound on the generation complexity for compositional problems:
$N(P,x_{1}...x_{k})\geq[\prod_{i=1}^{k}N(P,x_{i})]\cdot e^{\Delta\cdot\sum_{i=1}^{k}L_{i}}$ (4)
For some $\Delta>0$, where $L_{i}$ is the minimal length solution to $x_{i}$. In words, this result implies that when sampling a solution to sub-problems in one context, $y_{1}...y_{k}\sim P(\cdot|x_{1}...x_{k})$, the probability for them all to be correct is exponentially lower (in total code length) than if we sample them in parallel contexts: $y_{1}\sim P(\cdot|x_{1})...y_{k}\sim P(\cdot|x_{k})$. The implication of this result is that if we do not have access to tests for the intermediate steps $y_{1}...y_{k}$, but only for the final output $y_{1}\oplus...\oplus y_{k}$ (which is often the case as the model chooses how to decompose a problem), it is more beneficial to sample $y_{1}...y_{k}$ in parallel. This defines a compositional hardness of coding problems that originates from context processing limitations.
We note this result is not equivalent to showing a model’s success rate in compositional problems drops exponentially with the number of compositions, which was shown in Dziri et al. (2024), but rather it is a stronger result. Exponentially decreasing success rate in number of compositions means $N(P,x_{1}...x_{k})\geq\exp(c\cdot k)$ for some $c>0$. But the product of success rates of performing the subtasks in parallel $\prod_{i=1}^{k}N(P,x_{i})$ is naturally also exponential in $k$. Hence such results do not imply anything about whether it is more beneficial to sample in one context or to parallelize. In our results, we show their ratio is exponentially growing, $N(P,x_{1}...x_{k})/(\prod_{i=1}^{k}N(P,x_{i}))\geq e^{\Delta\cdot\sum_{i=1}^{k}L_{i}}$, which does reveal the advantage to sample in parallel.
@@ 3.2 Screening in Autoregressive Models
In the above, we have motivated to look at LLM’s abilities to solve multiple tasks in a single context compared with solving them in separate contexts. Here we lay theoretical grounds to explain the difference between the two scenarios based on the autoregressive nature of LLMs. For ease of presentation, below we present our assumptions for compositions of two problems, but note that it can inductively be increased to any number.
Typically, latent representations of the model contain information about the context beyond next token prediction, e.g. structure of the solution to problems (Ye et al., 2024). Thus when composing two code problems, we expect that during generation, the representations of the second program to contain information about the first and vice versa, which is grammatically similar (same programming language) but semantically very different. As a result, this additional information creates noise that can harm the generation process.
Formally, we denote by $r^{(L)}(x)$ the model’s last hidden layer representation of the prompt $x$, by $U$, the model’s unembedding matrix (hidden dimension to vocabulary). The logit of the $i$’th token is thus defined as the token’s score before the softmax operation: $\langle r^{(L)}(x),U^{T}e_{i}\rangle$, where $e_{i}$ is the one-hot vector of the token. The probability distribution at each decoding step is the softmax applied to the logits, $P_{LLM}(i|x)=softmax(\langle r^{(L)}(x),U^{T}e_{i}\rangle)$.
@@
Essentially, this means the model attempts to generate the same solutions as in the non-compositional case, but noise may interfere in the process. The projection of this noise onto the dictionary creates noise in the logits during decoding, which can lead the model to make mistakes. It is worth mentioning that while theoretically after generating $y_{1}$, it may serve as an in-context example for $y_{2}$, contradicting the noise argument, in practice when composing two complex problems that are semantically different, the model cannot learn from the first problem how to improve its success rate on the second. This remains true for problems decomposed with a chain of thoughts, in which each sub-task serves a different purpose and implementing it requires different tools, hence success in previous sub-tasks are not expected to create an in-context boot-strapping effect.
As the two problems $x_{1},x_{2}$ and their solutions $y_{1},y_{2}$ may be different semantically, we do not expect the noise to “push" the model towards correct solutions more than to incorrect solutions, so when projected onto the vocabulary, $V$, the noise on the logit of a correct next token, $i_{CNT}$, minus the noise on an incorrect token $i$, denoted, $\langle noise,U^{T}e_{i_{CNT}}\rangle-\langle noise,U^{T}e_{i}\rangle$, should be symmetric on average. Additionally it should be bounded within some range $[-M,+M]$, as it changes the hidden representation to a finite extent. In practice, we only expect this to be true for the high probability tokens, as the vocabulary $V$ is very large, and some low probability tokens may be systematically enhanced. To avoid this issue, we make our assumptions only on the weighted average of the noise, where the weights are given by the probability mass that the model assigns them. This way, low probability tokens with asymmetric noise or large norms receive low weight and are averaged with the noise of other tokens. Denote by $P(i|c)$ the probability assigned to the $i$’th token given the context $c$. We make our assumptions on the weighted average of the noise on incorrect token logits minus correct token logits:
$X=\frac{\sum_{i\in V\backslash\{i_{CNT}\}}P(i|c)\langle noise,U^{T}e_{i}-U^{T}e_{i_{CNT}}\rangle}{\sum_{i\in V\backslash\{i_{CNT}\}}P(i|c)}$ (7)